OpenStudy (anonymous):

1) Which of the following will form a right triangle? 7, 24, 25 7, 23, 24 7, 25, 26

OpenStudy (anonymous):

Pythagorean theorem

OpenStudy (dan815):

a^2+b^2=c^2 a and b are the smaller sides

OpenStudy (anonymous):

is the answer the first set of numbers?

OpenStudy (anonymous):

Ist.. use pythagorus.. a^2 +b^2=c^2

OpenStudy (anonymous):

49 + 576= 625

OpenStudy (anonymous):

25x25=625

OpenStudy (nincompoop):

pythagorean triples